Assume you opened and deposited $1000.00 into a savings account that pays 4% per annum.

  1. If the bank compounds interest annually, how much will you have in your account at the end of 3 years (assuming no deposits or withdrawals are made for 3 years)?
  2. Find the balance if the bank compounds interest quarterly under the same conditions.
  3. Find the balance if the bank compounds interest continuously under the same conditions.
Answer #1

Pg: - 0 A= P(1+ A: Future value P= Present Value Y: interest Rate to time period n: no- of times compound P: $ 1000 X= 4% n == 1000 (14 6.04413 = 1000 (1+0-01) = 1000% 1.12662 = 1126 82 After a year it will be - 11126.99 (0) Bornik. jotrest compound